# MAPO: Advancing Multilingual Reasoning through Multilingual-Alignment-as-Preference Optimization

### v2

@@ 3.1 Preference Estimation
We simply adopt a well-trained multilingual translation model to calculate the ‘alignment’ scores between answers in non-dominant and dominant languages as a preference. Since the translation model is optimized on a large scale parallel data in source and target languages $(\mathbf{Y},\mathbf{\bar{Y}})$ by maximizing conditional generation probability in Eq 3:
$\mathop{\arg\max}\limits_{\theta}P(\mathbf{\bar{Y}}|\mathbf{{Y}};\theta),$ (3)
with more aligned target language $\mathbf{\bar{Y}}$ to $\mathbf{Y}$, the higher the conditional generation probability $P(\mathbf{\bar{Y}}|Y)$. Then, we input the answers in non-dominant languages and force-decode a corresponding answer $\bar{Y}$ in dominant language for the ‘alignment’ score. Correspondingly, answers with higher ‘alignment’ scores in non-dominant languages are preferred during the following preference optimizations.
@@ 4.1 Experiment Setup22footnotemark: 2
Base Models: Our experiments include two base models. MathOctopus-7B is fine-tuned from LLaMA2-7B with the multilingual reasoning processes MGSM8KInstruct [^4] GSM8K translated to nine non-English languages. Chen et al. (2023). MetaMathOctopus-7B is built from MetaMath-7B Yu et al. (2023), which has stronger English reasoning abilities, by fine-tuning with MGSM8KInstruct as well.
@@
MathOctopus 13B 35.2 46.8 42.8 43.2 48.8 47.6 44.4 48.0 48.4 53.2 45.8
MAmmoTH ${}^{\dagger}$ 4.3 6.3 4.2 26.7 26.8 33.7 39.6 42.9 39.9 45.1 26.3
WizardMath ${}^{\dagger}$ 16.1 17.0 10.3 37.9 36.3 37.4 39.2 44.8 37.7 48.5 32.5
@@
Table 2: Model Performances on MSVAMP test set. “Avg” represents the average performance on ten languages and bold text denotes the best results within the same size. Results marked with ${}^{\dagger}$ come from Chen et al., 2023.
@@
  • ACR (Answer Consistency Ratio) : Let $m$ denote the set of items answered correctly in English, and $n$ denote those answered correctly in non-English. ACR is then calculated as: $\text{ACR}=\frac{|m\cap n|}{|n|}$. Higher ACR indicates a greater degree of overlap in reasoning capabilities between non-English and English languages.
@@
5 Experiment Results
@@
MathOctopus 13B 42.4 39.2 44.8 38.8 49.6 45.2 48.4 53.6 43.2 54.8 46.0
MAmmoTH ${}^{\dagger}$ 3.6 4.8 2.4 10.8 17.2 26.0 33.2 32.4 32.8 49.6 21.3
WizardMath ${}^{\dagger}$ 2.0 4.0 3.4 24.0 22.4 30.8 30.4 34.8 30.4 47.6 23.0
@@
Table 3: Model Performances on MGSM test set. “Avg” represents the average performance in ten languages and bold text denotes the best results within the same model size. Results marked with ${}^{\dagger}$ come from Chen et al., 2023.
Model Bn Th Sw Ja Zh Ru De Es Fr En Avg
GPT-3.5-Turbo 36.2 42.6 47.2 58.1 60.6 42.6 41.5 54.9 39.4 70.6 49.4
MathOctopus 13B 32.2 35.8 38.2 42.7 46.0 38.8 36.0 45.2 37.5 50.5 40.3
MAmmoTH 7B 1.7 4.7 4.5 23.5 29.9 26.9 34.7 37.9 36.5 42 24.2
WizardMath 7B 9.8 13 6.2 31.8 36 34.3 34.1 39.9 36.5 45.4 28.7
MetaMath 7B 13.2 19.2 12.1 50.7 53.3 52.9 54.2 56.5 56.7 63.3 43.2
MathOctopus 7B 26.6 30.9 34.3 40.9 44.4 36.0 32.6 42.0 36.2 46.9 37.1
+ m-RFT 38.0 42.9 41.8 48.2 51.6 45.2 42.9 49.3 42.9 51.2 45.4
+ MAPO-DPO(ours) 41.8 45.8 46.9 52.9 54.4 49.9 50.7 54.0 51.4 55.9 50.4
+ m-RFT 40.3 45.0 45.2 51.4 57.8 51.6 51.6 58.8 50.1 65.3 51.7
+ MAPO-DPO(ours) 40.7 46.0 45.0 58.2 59.3 53.1 51.4 57.4 52.0 66.1 52.9
@@ 5.1 Preference Optimization Improves Multilingual Reasoning Effectively
Experimental results in Table 2,3,4 consistently demonstrate that our method has effectively enhanced the reasoning capabilities of the base model and achieved state-of-the-art performance. More specifically, the improvement is most impressive on the out-of-domain dataset MSVAMP, where we achieved an average accuracy improvement of 16.2% and 11.7%, over the two base models, MathOctopus7B and MetaMathOctopus7B, respectively. We also observe significant improvements in MGSM and MNumGLUESub.
From the perspective of languages, most languages have improvement after alignment and it is more significant in some low-resource languages that previously underperformed. Our method has increased the accuracy on MSVAMP for Bengali, Thai, Swahili, and Japanese by 21.1%, 19.3%, 16.6%, and 18.7%, respectively.
@@ 6 Analysis
English Question
182 students suggested adding mashed potatoes while others suggested adding bacon to the menu . if 166 more students suggested adding bacon than those that suggested mashed potatoes how many students suggested bacon ?
Chinese Question
182名学生建议在菜单中加入土豆泥,而其他人建议加入培根。如果有比建议加入土豆泥的学生多166人建议加入培根,那么有多少学生建议加入培根?
English Solution
The number of students who suggested mashed potatoes is 182. The number of students who suggested bacon is 182+166=348.
Chinese Solution [Before Alignment]
建议加入土豆泥的学生比建议加入培根的学生多166人,所以两边减去166得到差值为182 - 166 = 16,因此,建议加入培根的学生人数为16。(There are 166 more students recommending mashed potatoes than bacon so subtract 166 from both sides to get a difference of 182-166 = 16. Therefore, the number of students recommended to join Bacon is 16.)
[PPL]: 2.65
[After Alignment]
建议加入土豆泥的学生人数是182。建议加入培根的学生人数是182 + 166 = 348。 (The number of students who suggested mashed potatoes is 182. The number of students who suggested bacon is 182+166=348.)
[PPL]: 0.97
@@ 6.4 Alignment Reasoning for Improvements
The example in Table 6 directly illustrates how our framework can improve reasoning ability by aligning reasoning in other languages to English. For the given English problem, the basic model correctly analyzes that an addition should be conducted to find the number of students recommended to join Bacon. However, when given the same problem in Chinese, LLM misjudges the relationship of variables and writes a significantly different reasoning process. This directly proves that although the baseline model has been fine-tuned with multilingual reasoning data, its reasoning and thoughts are inconsistent, whereas non-English reasoning is more prone to errors. After alignment, the reasoning thoughts are more similar to the English answer, and the reasoning thoughts are also corrected.
@@ 7 Conclusion
Similar to previous work on preference optimization, our method necessitates a policy model that has the preliminary multilingual reasoning capabilities through SFT. Meanwhile, due to limitations in computational resources, our experiments are confined to aligning models of 7B size and exploring two preference optimization algorithms, PPO and DPO. Should resources permit, we aim to extend our exploration to models of 13B and 70B sizes and examine the performance of a broader spectrum of preference optimization algorithms.
@@ Appendix A Experiment Details
Our code is primarily based on the trl 666 https://github.com/huggingface/trl , with some minor modifications made. The modified code will also be made available at our project.

### v3

@@ 3.1 Preference Estimation
We simply adopt a well-trained multilingual translation model to calculate the ‘alignment’ scores between answers in non-dominant and dominant languages as a preference. Since the translation model is optimized on a large scale parallel data in source and target languages $(Y,\bar{Y})$ by maximizing conditional generation probability in Eq 3:
$\mathop{\arg\max}\limits_{\theta}P(\bar{Y}|{Y};\theta),$ (3)
with more aligned target language $\bar{Y}$ to $Y$, the higher the conditional generation probability $P(\bar{Y}|Y)$. Then, we input the answers in non-dominant languages and force-decode a corresponding answer $\bar{Y}$ in dominant language for the ‘alignment’ score. Correspondingly, answers with higher ‘alignment’ scores in non-dominant languages are preferred during the following preference optimizations.
@@ 4.1 Experiment Setup22footnotemark: 2
Base Models: Our experiments include four base models. MathOctopus-7/13B is fine-tuned from LLaMA2-7/13B with the multilingual reasoning processes MGSM8KInstruct [^4] GSM8K translated to nine non-English languages. Chen et al. (2023). Yu et al., 2023 proposed MetaMath-7/13B and MetaMath-Mistral-7B, which have stronger English reasoning abilities. We fine-tune these models with MGSM8KInstruct and get MetaMathOctopus-7/13B and MistralMathOctopus-7B.
@@
MAmmoTH 7B 4.3 6.3 4.2 26.7 26.8 33.7 39.6 42.9 39.9 45.1 26.3
WizardMath 7B 16.1 17.0 10.3 37.9 36.3 37.4 39.2 44.8 37.7 48.5 32.5
@@
MistralMathOctopus 7B 47.0 52.6 54.4 58.8 60.2 59.8 62.1 60.8 60.4 74.4 59.0
+ m-RFT 57.4 63.5 60.8 65.2 69.8 67.4 67.2 69.1 68.0 71.3 66.0
+ MAPO-DPO(ours) 62.9 71.3 71.4 73.7 76.0 74.9 77.8 78.1 79.0 81.1 74.6
MAmmoTH 13B 5.0 13.7 12.9 42.2 47.7 50.7 52.3 53.9 53.8 53.4 38.6
WizardMath 13B 13.7 16.3 12.5 29.5 37.0 43.8 48.7 49.4 49.4 56.3 35.7
MetaMath 13B 14.6 15.7 17.4 57.0 56.6 63.7 67.3 65.9 64.7 67.7 49.1
MathOctopus 13B 35.2 46.8 42.8 43.2 48.8 47.6 44.4 48.0 48.4 53.2 45.8
+ m-RFT 43.4 50 52.1 54.9 55.4 57.1 59.2 56.4 59.5 55.2 54.3
+ MAPO-DPO(ours) 51.8 58.9 56.4 60.4 58.8 62.1 63.5 62.0 61.7 65.0 60.1
MetaMathOctopus 13B 41.6 52.1 50.9 57.3 53.1 59.1 60.1 61.1 60.8 66.8 56.3
+ m-RFT 48.1 59.6 61.4 60.5 58.9 61.0 62.7 65.3 64.3 65.4 60.7
+ MAPO-DPO(ours) 54.7 64.7 62.9 69.0 68.2 68.2 69.5 70.6 71.3 70.5 67.0
Table 2: Model Performances on MSVAMP test set. “Avg” represents the average performance on ten languages and bold text denotes the best results within the same size. Results marked with come from Chen et al., 2023.
@@
  • ACR (Answer Consistency Ratio) : Let $m$ denote the set of math questions answered correctly in English, and $n$ denote those answered correctly in non-English. ACR is then calculated as: $\text{ACR}=\frac{|m\cap n|}{|n|}$. Higher ACR indicates a greater degree of overlap in reasoning capabilities between non-English and English languages.
@@
5 Experiment Results66footnotemark: 6
77footnotetext: To facilitate further research, we launch a multilingual reasoning leaderboard: https://huggingface.co/spaces/kevinpro/Open-Multilingual-Reasoning-Leaderboard
@@ 5 Experiment Results66footnotemark: 6
MAmmoTH 7B 3.6 4.8 2.4 10.8 17.2 26.0 33.2 32.4 32.8 49.6 21.3
WizardMath 7B 2.0 4.0 3.4 24.0 22.4 30.8 30.4 34.8 30.4 47.6 23.0
@@
MistralMathOctopus 7B 44.0 54.4 50.4 55.6 59.2 58.8 62.4 62.0 56.8 76.0 58.0
+ m-RFT 41.2 46.8 46.8 48.4 57.2 62.8 61.6 59.2 57.6 72.0 55.4
+ MAPO-DPO(ours) 55.2 60.4 61.6 58.0 74.0 70.8 67.6 74.0 69.2 82.0 67.3
MAmmoTH 13B 3.6 5.2 1.6 19.2 31.2 36.8 45.6 50.0 39.6 56.4 28.9
WizardMath 13B 6.4 5.6 5.6 22.0 28.0 34.4 40.4 45.6 42.0 52.8 28.3
MetaMath 13B 11.6 6.4 7.6 42.8 49.2 63.6 64.8 65.2 65.2 67.2 44.4
MathOctopus 13B 42.4 39.2 44.8 38.8 49.6 45.2 48.4 53.6 43.2 54.8 46.0
+ m-RFT 29.6 30.8 34.4 36.4 40.4 39.2 42.0 42.8 40.4 48.0 38.4
+ MAPO-DPO(ours) 38.8 46.8 42.0 47.6 53.6 49.2 52.0 54.4 46.4 54.0 48.5
MetaMathOctopus 13B 34.4 42.8 41.6 49.2 52.8 54.4 54.4 59.2 53.6 71.6 51.4
+ m-RFT 22.8 29.6 30.4 35.2 39.2 40.0 43.6 43.6 41.2 59.2 38.5
+ MAPO-DPO(ours) 44.8 47.6 55.2 56.0 59.6 59.2 59.2 63.6 62.8 71.6 58.0
Table 3: Model Performances on MGSM test set. “Avg” represents the average performance in ten languages and bold text denotes the best results within the same model size. Results marked with come from Chen et al., 2023.
@@ 5.1 Preference Optimization Improves Multilingual Reasoning Effectively
Experimental results in Table 2,3,7 consistently demonstrate that our method has effectively enhanced the reasoning capabilities of the various base models and achieved state-of-the-art performance. More specifically, the improvement is most impressive on the out-of-domain dataset MSVAMP, where we achieved an average accuracy improvement of 16.2% and 14.7% MathOctopus7B and MathOctopus13B, respectively. Even for the most powerful 7B model MistralMathOctopus 7B, our method can further boost its performance to an impressive 74.6%. We also observe significant improvements in MGSM and MNumGLUESub.
From the perspective of languages, most languages have improvement after alignment and it is more significant in some low-resource languages that previously under-performed. Take MathOctopus 7B as an example, our method has increased the accuracy on MSVAMP for Bengali, Thai, Swahili, and Japanese by 21.1%, 19.3%, 16.6%, and 18.7%, respectively.
@@ 6.1 Preference Estimating on a Dataset Different from SFT
System MSVAMP MGSM
MetaMathOctopus7B 53.0 45.5
+ Ours & NLLB 600M 61.1 48.9
+ Ours & MBART-MMT-600M 59.9 49.6
+ Ours & M2M-1.2B 61.4 49.3
Table 5: Average accuracy on ten languages on MSVAMP and MGSM. Note that we only conduct one iteration MAPO-DPO with different translation models.
English Question 182 students suggested adding mashed potatoes while others suggested adding bacon to the menu . if 166 more students suggested adding bacon than those that suggested mashed potatoes how many students suggested bacon ?
Chinese Question {CJK*}UTF8gkai 182名学生建议在菜单中加入土豆泥,而其他人建议加入培根。如果有比建议加入土豆泥的学生多166人建议加入培根,那么有多少学生建议加入培根?
English Solution The number of students who suggested mashed potatoes is 182. The number of students who suggested bacon is 182+166=348.
Chinese Solution [Before Alignment]
{CJK*}UTF8gkai 建议加入土豆泥的学生比建议加入培根的学生多166人,所以两边减去166得到差值为182 - 166 = 16,因此,建议加入培根的学生人数为16。(There are 166 more students recommending mashed potatoes than bacon so subtract 166 from both sides to get a difference of 182-166 = 16. Therefore, the number of students recommended to join Bacon is 16.)
{CJK*}UTF8gkai [PPL]: 2.65
[After Alignment]
{CJK*}UTF8gkai 建议加入土豆泥的学生人数是182。建议加入培根的学生人数是182 + 166 = 348。 (The number of students who suggested mashed potatoes is 182. The number of students who suggested bacon is 182+166=348.)
{CJK*}UTF8gkai [PPL]: 0.97
7 Robustness over Different Translation Models
To verify the robustness of our framework across various translation models, we conduct experiments using MBART-MMT-600M  [^8] https://huggingface.co/facebook/mbart-large-50-many-to-many-mmt and M2M-1.2B  [^9] https://huggingface.co/facebook/m2m100_1.2B, which have the different model architectures, scales, and training datasets. The results, presented in Table 5, demonstrate that our framework consistently achieves performance improvements. The performance is even slightly enhanced when with M2M-1.2B. These experimental results conclusively confirm the robustness of our framework. In addition, we have conducted experiments on translation models of different scales, and the results also demonstrate the robustness of our method. For more details, please refer to the Appendix D.
@@ 7.2 Alignment Reasoning for Improvements
The example in Table 6.1 directly illustrates how our framework can improve reasoning ability by aligning reasoning in other languages to English. For the given English problem, the basic model correctly analyzes that an addition should be conducted to find the number of students recommended to join Bacon. However, when given the same problem in Chinese, LLM misjudges the relationship of variables and writes a significantly different reasoning process. This directly proves that although the baseline model has been fine-tuned with multilingual reasoning data, its reasoning and thoughts are inconsistent, whereas non-English reasoning is more prone to errors. After alignment, the reasoning thoughts are more similar to the English answer, and the reasoning thoughts are also corrected.
@@ 8 Conclusion
Similar to previous work on preference optimization, our method necessitates a policy model that has the preliminary multilingual reasoning capabilities through SFT. Meanwhile, due to limitations in computational resources, our experiments are confined to aligning models of 7B and 13B size and exploring two preference optimization algorithms, PPO and DPO. Should resources permit, we aim to extend our exploration to models of 70B sizes and examine the performance of a broader spectrum of preference optimization algorithms.
@@ Appendix A Experiment Details
Our code is primarily based on the trl 101010 https://github.com/huggingface/trl , with some minor modifications made. The modified code will also be made available at our project.
@@ Appendix C Used Scientific Artifacts
Appendix D Scaling of Translation Model
Model Bn Th Sw Ja Zh Ru De Es Fr En Avg
GPT-3.5-Turbo 36.2 42.6 47.2 58.1 60.6 42.6 41.5 54.9 39.4 70.6 49.4
MAmmoTH 7B 1.7 4.7 4.5 23.5 29.9 26.9 34.7 37.9 36.5 42 24.2
WizardMath 7B 9.8 13 6.2 31.8 36 34.3 34.1 39.9 36.5 45.4 28.7
MetaMath 7B 13.2 19.2 12.1 50.7 53.3 52.9 54.2 56.5 56.7 63.3 43.2
MathOctopus 7B 26.6 30.9 34.3 40.9 44.4 36.0 32.6 42.0 36.2 46.9 37.1
+ m-RFT 38.0 42.9 41.8 48.2 51.6 45.2 42.9 49.3 42.9 51.2 45.4
+ MAPO-DPO(ours) 41.8 45.8 46.9 52.9 54.4 49.9 50.7 54.0 51.4 55.9 50.4
+ m-RFT 40.3 45.0 45.2 51.4 57.8 51.6 51.6 58.8 50.1 65.3 51.7
+ MAPO-DPO(ours) 40.7 46.0 45.0 58.2 59.3 53.1 51.4 57.4 52.0 66.1 52.9
MistralMathOctopus 7B 45.2 50.7 49.5 56.5 65.3 59.1 51.4 62.1 53.9 74.6 56.8
+ m-RFT 52.7 60.6 61.0 67.2 70.8 65.9 61.2 71.6 64.4 78.3 65.4
+ MAPO-DPO(ours) 62.3 64.6 61.6 72.1 75.1 68.0 69.3 74.4 74.2 78.5 70.0
MAmmoTH 13B 6.8 10.5 10.4 31.6 38 41.1 41.4 43.3 42.6 46.7 31.2
WizardMath 13B 9.8 14.3 12.4 30.5 39.0 36.5 35.2 43.9 39.2 47.8 30.9
MetaMath 13B 10.9 16.0 16.0 55.2 57.4 56.5 58.9 60.6 58.0 64.2 45.4
MathOctopus 13B 42.4 39.2 44.8 38.8 49.6 45.2 48.4 53.6 43.2 54.8 46.0
+ m-RFT 44.1 49.9 51.0 51.0 55.6 49.0 47.1 53.3 46.1 57.4 50.5
+ MAPO-DPO(ours) 49.0 53.3 52.2 55.0 57.4 53.3 52.2 55.9 50.7 58.8 53.8
MetaMathOctopus 13B 34.4 42.8 41.6 49.2 52.8 54.4 54.4 59.2 53.6 71.6 51.4
+ m-RFT 40.1 51.6 47.5 60.5 62.0 58.9 54.8 62.5 54.2 66.5 55.9
+ MAPO-DPO(ours) 52.9 55.4 55.0 67.2 65.0 54.8 54.4 65.2 57.4 70.4 59.8